A hardware production company produces a “Hex Bolt”. The specifications for this bolt are very strict. The specified diameter is 4.770 mm with an allowed tolerance is +/ − 0.001mm.
Currently, the process is centered, with a standard deviation of 0.0001 mm.

What is the CPK value of this process?

Select one:

a. 0.002

b. 1.67

c. 3.33

d. 6.67

e. None of the above are correct
Answer #1

c. 3.33

CpK = Min(CpU,CpL)

CpU = (Upper Specification Limit - Mean)/3*Standard Deviation

= (4.771-4.77)/3*0.0001

= 3.33

CpL = (Mean - Lower Specification Limit)/3*Standard Deviation

= (4.77-4.769)/3*0.0001

= 3.33

CpK = Min(3.33,3.33)

= 3.33